Given two goods named Omega & Alpha, with Omega along the vertical axis, and Alpha on the
horizontal. Assume that the budget is constant at 100 pesos, the price of good Omega is fixed at 5
pesos. If the Price of Alpha changes from 1 peso to 2 pesos, how many units of Omega can be
bought in combination with 50 pieces of Alpha at its initial price ?
1
Expert's answer
2021-09-22T08:56:21-0400

Solution:

Budget constraint: I = PxX + PyY

Where: I = Budget income = 100

           X = Alpha

           Y = Omega

          Px = Price of Alpha

          Py = Price of Omega = 5

The initial price of Alpha = 1 peso

100 = X + 5Y

When 50 pcs of Alpha are purchased:

100 = 50 + 5Y

100 – 50 = 5Y

50 = 5Y

Y = 10

Omega = 10 units


The units of Omega that can be bought in combination with 50 pieces of Alpha at its initial price = 10 units